OpenStudy (anonymous):

A mirror with a parabolic cross section is used to collect sunlight on a pipe located at the focus of the mirror. The pipe is located 2 inches from the vertex of the mirror. Write an equation of the parabola that models the cross section of the mirror. Assume that the parabola opens upward.

OpenStudy (anonymous):

@sami-21 & @satellite73

OpenStudy (anonymous):

Help ? Anyone ?

OpenStudy (anonymous):

assuming the vertex of the parabola is at the origin, the equation will look like this: \(\large 4py=x^2 \)

OpenStudy (anonymous):

Okay so what do i do ?

OpenStudy (anonymous):

since the distance from the vertex to the focus is 2, then p=2 so your equation is : \(\large 4(2)y=x^2 \) or simply \(\large 8y=x^2 \) or \(\large y=\frac{1}{8}x^2 \)

OpenStudy (anonymous):

It would be postive right ?

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

yw :)